OpenStudy (anonymous):

identify the pattern 45 38 31 24 ( ) ( ) what gose in the blank

jimthompson5910 (jim_thompson5910):

Look at the difference between each value. What is that difference?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

- 7

jimthompson5910 (jim_thompson5910):

good, so you're subtracting 7 from each term so the next term is then....?

OpenStudy (anonymous):

i got it wats 50 40 43 33 ( ) 26 ( ) fill in the blank

jimthompson5910 (jim_thompson5910):

change from 50 to 40 is -10 (since 40-50 = -10). So you're subtracting 10 but the change from 40 to 43 is +3 (since 43-40 = 3). So you're adding 3 and then notice the change comes back to -10 since 33-43 = -10 This pattern continues and it is -10 followed by +3 So after 33, we add 3 to get 33+3 = 36 Then we subtract 10 to get 36-10 = 26 Finally, we add 3 to get 26+3 = 29